Suppose that Mountain Star Bank discovers that its reserves will temporarily fall slightly below those legally...

  1. Suppose that Mountain Star Bank discovers that its reserves will temporarily fall slightly below those legally required. How might it temporarily remedy this situation through the Federal funds market? Now assume Mountain Star finds that its reserves will be substantially and permanently deficient. What remedy is available to this bank?
  2. How would a decrease in the reserve requirement affect the (a) size of the money multiplier, (b) amount of excess reserves in the banking system, and (c) extent to which the system could expand the money supply through the creation of checkable deposits via loans?

Show your work for the problems below. Explain and show how the results were determined for full credit.

3. Suppose that Serendipity Bank has excess reserves of $8,000 and checkable deposits of $150,000. If the reserve ratio is 20 percent, what is the size of the bank’s actual reserves?  

4. The Third National Bank has reserves of $20,000 and checkable deposits of $100,000. The reserve ratio is 20 percent. Households deposit $5000 in currency into the bank and that currency is added to reserves. What level of excess reserves does the bank now have?
Answer #1

1. The federal funds market is market where banks can borrow funds for overnight to meet their reserve requirements. The mountain start bank can use the federal funds market to borrow the funds from other banks or federal reserve to meet the requirements. But this is only for temporary reserve deficiency.

If the reserves are deficient permanently, then the bank will have to raise money for longer term. This means getting more deposits (through offering more attractive rates etc) or raisig money through other method such as equity.

2. The moey multiplier is given by

1/reserve ratio.

So when the reserve requirement decreases, money multiplier increases.

It also increases excess reserves as banks were earlier holding higher reserves (because the required ratio was higher).

As the multiplier increases, so does the capacity of expanding money through creation of loans.

3. The 20% reserve ratio requirement means that the bank needs to hold

150000*20%=30000 in reserves.

As the bank holds 8000 in excess, it means its total reserves are

30000+8000=38000

4. The bank earlier had 100000 in deposits. Because the reserve requirement was 20%, it means that the bank needed to hold $20000. So earlier there were no extra reserves.

Now the total deposits are 105000.

Required reserves=105000*.2=21000.

But actually the bank is holding 20000+5000=25000 in reserves now.

So, excess reserves=25000-21000=4000.